# Sub-Riemannian Geometry, Mixing, and the Holonomy of Optimal Mass Transport

### v1

@@ 4.3 The Isoparallel Mass Transport Problem
Problem 1 (Isoparallel Mass Transport (IMT))
@@
Evidently, the essence of Problem 1 is to effect transport that links two distributions on the base space $\mathcal{N}(\mathbb{R}^{n})$, corresponding to $\Sigma_{\rm in}$ and $\Sigma_{\rm fin}$, respectively, ensuring a specified terminal positioning of the particles while minimizing the quadratic cost that is associated with the length that is being traversed under the control of potential forces.
The IMT problem is an instance of Montgomery’s isoparallel problem [18] expressed in the context of OMT. It is equivalent to the problem of finding a geodesic with respect to the OMT sub-Riemannian structure. Unlike the general case of the latter however, sub-Riemannian geodesics associated to solutions of the isoparallel problem are never abnormal. For a proof of this fact, we refer the reader to [18, Theorem 1]. In fact, Proposition 6 already rules out the existence of nontrivial abnormal geodesics for the OMT sub-Riemannian structure, which implies that solutions to the IMT problem, when they exist, are necessarily projections of normal sub-Riemannian geodesics.
@@
Proof: Fixing $\Phi_{\text{in}}\in\pi^{-1}(\Sigma_{\text{in}})$, the IMT problem is equivalent to the problem of finding the normal sub-Riemannian geodesic between $\Phi_{\text{in}}$ and $\Phi_{\text{des}}\Phi_{\text{in}}$. From Proposition 7, there must exist a unique smooth curve $\Lambda_{.}:[0,1]\rightarrow\text{GL}^{+}(n)$ such that the sought-after geodesic satisfies
@@
The proof is concluded by observing that $\Pi_{t}=\dot{\Sigma}_{t}^{\star}$. $\Box$
@@
Sub-Riemannian geodesics between points that are sufficiently close, with respect to the sub-Riemannian distance $d_{SR}$, are guaranteed to exist [19, Theorem 1.18]. The standard sufficient condition, see [18, Proposition 1], to guarantee global existence of sub-Riemannian geodesics, and consequently, global existence of solutions to the isoparallel problem, include the condition that the Riemannian structure on the base space of the principal bundle is complete, i.e. that $\text{Sym}^{+}(n)$ equipped with the Wasserstein-Otto metric (40), abbreviated as $(\text{Sym}^{+}(n),\bar{\mathcal{G}})$ for short, is a complete Riemannian manifold. Unfortunately, it is well-known that $(\text{Sym}^{+}(n),\bar{\mathcal{G}})$ is not complete. As shown in Remark 1, nontrivial geodesics in $(\text{Sym}^{+}(n),\bar{\mathcal{G}})$ may be extended in one but never in both directions. This alone is sufficient to prove that the sub-Riemannian structure of OMT is not geodesically complete. Indeed, we have the following proposition which follows from the fact that the OMT sub-Riemannian structure is of the bundle-type [19, Chapter 11].
Proposition 11
@@ 4.4 Examples: sub-Riemannian geodesics
$\pi^{-1}(\Sigma)$ =$=\{\Sigma^{\frac{1}{2}}\Theta~{}|~{}\Theta\in\text{SO}(2)\},$ for all $\Sigma\in\text{Sym}^{+}(n)$.
@@
If $\Phi_{\text{des}}=\Phi^{\star}$ is the matrix associated with the Monge map, i.e. the matrix given by
@@
As can be observed, the tracer particles travel along straight lines (geodesics of $\mathbb{R}^{n}$) which is a characteristic feature of McCann geodesics.
@@
If $\Sigma_{\text{fn}}=\Sigma_{\text{in}}=I$, then $\Phi_{\text{des}}=\Theta\in\text{SO}(2)$, and a sub-Riemannian geodesic realizing the parallel transport map $\Phi_{\text{des}}$ will necessarily connect two points of the same fiber $\pi^{-1}(I)=\text{SO}(2)$. In fact, there’s an entire 1-parameter family of sub-Riemannian geodesics that realize $\Phi_{\text{des}}$ for every $\Phi_{\text{des}}\in\text{SO}(2)$ sufficiently close to $I$! Such family has a simple characterization in terms of any of its elements. Namely, if $\Sigma_{.}$ is a solution of the IMT problem with the above data, then $\tilde{\Theta}\sharp\Sigma_{.}$ is also a solution for every $\tilde{\Theta}\in\text{SO}(2)$. This phenomenon is ubiquitous in sub-Riemannian geometry [19, Section 1.8]. A sample trajectory, with $\Theta$ as prescribed in (59), is shown in Fig. 5. As can be observed, the tracer particles have rotated an angle of $\approx 28.6^{\circ}$ with respect to the origin after traversing the curve once.
@@ 5 Application: Transportation Cycles
In contrast to the formalism of uncertainty control where the focus is on regulating distributions, see e.g., [28, 29], herein the control protocol needs to guide individual particles along closed orbits. To this end, it is of essence that their position is “registered” relative to one another and relative to their common distribution. In other words, before designing a control law, one needs to know in advance how the particles are to be distributed when they arrive at the $k$th point of the cycle.
Figure 5: Trajectories of three tracer particle traversing the isoholonomic curve connecting $\Sigma_{\text{in}}=I$, $\Sigma_{\text{fn}}=I$, with the holonomy prescribed in (59). The dashed black ellipses mark intermediate covariances along the trajectory.
@@
$\triangle^{i,j}_{t}$ :=$:=\begin{cases}(I+3t(\Phi^{\star}_{\text{ref}\rightarrow i}-I)){\sharp}{\Sigma}_{\text{ref}},&t\in\left[0,\frac{1}{3}\right),\\ \Sigma^{i\mapsto j}_{\tau},\text{ with $\tau=3t-1$,}&t\in\left[\frac{1}{3},\frac{2}{3}\right),\\ (I+(3t-2)(\Phi^{\star}_{j\rightarrow\text{ref}}-I)){\sharp}\mathbf{\Sigma}_{j},&t\in\left[\frac{2}{3},1\right].\\ \end{cases}$
@@
In particular, $\Sigma^{i\mapsto j}_{.}$ is the curve traversed by the distribution of the particles in their flight from $\mathbf{\Sigma}_{i}$ to $\mathbf{\Sigma}_{j}$. From (61), it is clear that .$\triangle^{i,j}_{.}$ is a closed curve in $\text{Sym}^{+}(n)$. As can be verified, the holonomy of .$\triangle^{i,j}_{.}$ is
@@
$\Sigma_{t}$ :=$:=\begin{cases}\Sigma^{1\mapsto 2}_{\tau},\text{ with $\tau=Nt$,}&t\in\left[0,\frac{1}{N}\right],\\ \Sigma^{2\mapsto 3}_{\tau},\text{ with $\tau=Nt-1$,}&t\in\left[\frac{1}{N},\frac{2}{N}\right],\\ ~{}~{}\vdots~{}~{}&\hfill\vdots\hfill\\ \Sigma^{N\mapsto 1}_{\tau},\text{ with $\tau=N(t-1)+1$,}&t\in\left[\frac{N-1}{N},1\right].\\ \end{cases}$ (63)
@@
$1\rightarrow 2\rightarrow\dots\rightarrow N\rightarrow 1$
@@
Proof: By construction of the curve $\Sigma_{.}$, we have
@@
The proposition follows by expanding the product and sequentially applying the identity $\Phi^{\star}_{\text{ref}\rightarrow i}\Phi^{\star}_{i\rightarrow\text{ref}}=I.$ An alternative (geometric) proof for the simple case $N=3$ is sketched in Figure 8. The case of arbitrary $N$ is similar. $\Box$
@@
A schematic that compares the solution to the above problem with the McCann “edge” $(\mathbf{\Sigma}_{i},\mathbf{\Sigma}_{j})$ is shown in Figure 8. Returning to the Wasserstein polygon with vertices $\{\mathbf{\Sigma}_{1},\mathbf{\Sigma}_{2},\dots,\mathbf{\Sigma}_{N}\}$, it is evident that the choice of $\Sigma_{\rm ref}$ dictates the sub-Riemannian length of the edges, and hence the sub-Riemannian “perimeter” of the polygon, which represents the control cost in steering the ensemble along the cycle, as specified. Below we explore and compare two different options for choosing $\Sigma_{\text{ref}}$. Optimization of the control cost over $\Sigma_{\rm ref}$ will be tackled in future work.
@@ 6 Epilogue
The impact of Monge’s problem, to establish an optimal correspondence between distributions, has been foundational across mathematics, engineering, and the sciences [2, 32]. It constitutes a basic paradigm in stochastic control [28, 5] and has provided a setting for addressing a host of problems in statistics, probability theory, thermodynamics, economics, machine learning, to name a few.
Our motivation in this work has been to explore and bring to life a hitherto unstudied angle of a Lagrangian viewpoint, where the actual trajectories and mixing of particles are of interest. Thus, our central theme has been to control Gaussian distributions using potential forces, and to characterize the holonomy of particle trajectories. The OMT principal bundle structure duly encodes particle trajectories by recording state transition matrices $\Phi_{t}$. Equivalently, the trajectories of $n$ representative tracer particles is encoded in the columns of $\Phi_{t}$, and these provide information on the holonomy of the flow.
Our main contribution has been to introduce and study the holonomy of the Monge-Kantorovich optimal mass transport. The key insight has been to encode the Eulerian constraint on the velocity field to be potential into the OMT principal Ehresmann connection. We then briefly explored the sub-Riemannian structure induced by this connection along with the Wasserstein-Otto metric in the context of Gaussian distributions. A more comprehensive study of this structure in finite and infinite dimensions is of interest and will be pursued.
A similar framework should prove useful in ensemble control [34, 35, 36], where trajectories of multi-species populations indexed by $s$, for instance, are encoded in collections of transition matrices, $\Phi^{s}_{t}$’s, on a suitable bundle $\pi$ projecting onto the configuration in a base space, e.g., of covariance data of the form $\Sigma_{t}=\sum_{s}\Phi^{s}_{t}(\Phi^{s}_{t})^{\intercal}$. Other variants of practical interest include modeling the flow of particles based on aggregate information and the location of tracer particles used to estimate the holonomy of the collection, and likewise, estimation of trajectories of dynamical systems driven by stochastic excitation based on sample trajectories. These will be the subject of forthcoming work.

### v2

@@ 4.3 The Isoparallel Mass Transport Problem
Problem 1 (Isoparallel Mass Transport (IMT)).
@@
[…]
The IMT problem is an instance of Montgomery’s isoparallel problem [31] expressed in the context of OMT. It is equivalent to the problem of finding a geodesic with respect to the OMT sub-Riemannian structure.
@@
Proof.
Fixing $\Phi_{\text{in}}\in\pi^{-1}(\Sigma_{\text{in}})$, the IMT problem is equivalent to the problem of finding the normal sub-Riemannian geodesic between $\Phi_{\text{in}}$ and $\Phi_{\text{des}}\Phi_{\text{in}}$. From Proposition 7, there must exist a unique smooth curve $\Lambda_{.}:[0,1]\rightarrow\text{GL}^{+}(n)$ such that the sought-after geodesic satisfies
@@
The proof is concluded by observing that $\Pi_{t}=\dot{\Sigma}_{t}^{\star}$. ∎
@@
Sub-Riemannian geodesics between points that are sufficiently close, with respect to the sub-Riemannian distance $d_{SR}$, are guaranteed to exist [32, Theorem 1.18]. The standard sufficient condition, see [31, Proposition 1], to guarantee global existence of sub-Riemannian geodesics, and consequently, global existence of solutions to the isoparallel problem, include the condition that the Riemannian structure on the base space of the principal bundle is complete, i.e. that $\text{Sym}^{+}(n)$ equipped with the Wasserstein-Otto metric (40), abbreviated as $(\text{Sym}^{+}(n),\bar{\mathcal{G}})$ for short, is a complete Riemannian manifold. Unfortunately, it is well-known that $(\text{Sym}^{+}(n),\bar{\mathcal{G}})$ is not complete. As shown in Remark 1, nontrivial geodesics in $(\text{Sym}^{+}(n),\bar{\mathcal{G}})$ may be extended in one but never in both directions. This alone is sufficient to prove that the sub-Riemannian structure of OMT is not geodesically complete. Indeed, we have the following proposition which follows from the fact that the OMT sub-Riemannian structure is of the bundle-type [32, Chapter 11].
Proposition 11.
@@ 4.4 Examples: sub-Riemannian geodesics
$\pi^{-1}(\Sigma)$ $=\{\Sigma^{\frac{1}{2}}\Theta~|~\Theta\in\text{SO}(2)\},$ for all $\Sigma\in\text{Sym}^{+}(n)$.
@@
[…]
black If $\Phi_{\text{des}}=\Phi^{\star}$ is the matrix associated with the Monge map, i.e. the matrix given by
@@
The ellipses in Fig. 5, as well as in Fig. 5 and Fig. 5, mark the evolution of the level set of the gaussian distribution to which the indicated tracer particles belong. As can be observed, the tracer particles travel along straight lines (geodesics of $\mathbb{R}^{n}$) which is a characteristic feature of McCann geodesics.
@@
If $\Sigma_{\text{fn}}=\Sigma_{\text{in}}=I$, then $\Phi_{\text{des}}=\Theta\in\text{SO}(2)$, and a sub-Riemannian geodesic realizing the parallel transport map $\Phi_{\text{des}}$ will necessarily connect two points of the same fiber $\pi^{-1}(I)=\text{SO}(2)$. In fact, there’s an entire 1-parameter family of sub-Riemannian geodesics that realize $\Phi_{\text{des}}$ for every $\Phi_{\text{des}}\in\text{SO}(2)$ sufficiently close to II! Such family has a simple characterization in terms of any of its elements. Namely, if $\Sigma_{.}$ is a solution of the IMT problem with the above data, then $\tilde{\Theta}\sharp\Sigma_{.}$ is also a solution for every $\tilde{\Theta}\in\text{SO}(2)$. This phenomenon is ubiquitous in sub-Riemannian geometry [32, Section 1.8]. A sample trajectory, with $\Theta$ as prescribed in (59), is shown in Fig. 5. As can be observed, the tracer particles have rotated an angle of $\approx 28.6^{\circ}$ with respect to the origin after traversing the curve once.
@@ 5 Application: Transportation Cycles
In contrast to the formalism of uncertainty control where the focus is on regulating distributions, herein the control protocol needs to guide individual particles along closed orbits. To this end, it is of essence that their position is “registered” relative to one another and relative to their common distribution. In other words, before designing a control law, one needs to know in advance how the particles are to be distributed when they arrive at the $k$th point of the cycle.
@@
[…]
@@
In particular, $\Sigma^{i\mapsto j}_{.}$ is the curve traversed by the distribution of the particles in their flight from $\mathbf{\Sigma}_{i}$ to $\mathbf{\Sigma}_{j}$. From (61), it is clear that $\triangle^{i,j}_{.}$ is a closed curve in $\text{Sym}^{+}(n)$. As can be verified, the holonomy of $\triangle^{i,j}_{.}$ is
@@
[…]
@@
$1\rightarrow 2\rightarrow\dots\rightarrow N\rightarrow 1.$
@@
Proof.
By construction of the curve $\Sigma_{.}$, we have
@@
The proposition follows by expanding the product and sequentially applying the identity $\Phi^{\star}_{\text{ref}\rightarrow i}\Phi^{\star}_{i\rightarrow\text{ref}}=I.$ An alternative (geometric) proof for the simple case $N=3$ is sketched in Figure 8. The case of arbitrary $N$ is similar. ∎
@@
A schematic that compares the solution to the above problem with the McCann “edge” $(\mathbf{\Sigma}_{i},\mathbf{\Sigma}_{j})$ is shown in Figure 8.
[…]
Remark 4.
In principle, it is possible to seek a horizontal curve of minimal length that visits the given distributions, in the specified order, which leads to a sequence of optimal control problems that are nonlinearly coupled through their boundaries. This problem is numerically demanding, and the approach of registering the given distributions with respect to a reference decouples the problem into simpler ones. $\Box$
@@
Remark 5.
The path along a cycle may be optimized over the parallel transport maps along each segment, subject to the condition of a trivial holonomy. The more restrictive scheme suggested in the current section to achieve a trivial holonomy, i.e., registering the distributions with respect to a reference, simplifies the computation of parallel transport maps. $\Box$
@@ 6 Epilogue
The impact of Monge’s problem, to establish an optimal correspondence between distributions, has been foundational across mathematics, engineering, and the sciences [44, 36]. It constitutes a basic paradigm in stochastic control [9, 10] and has provided a setting for addressing a host of problems in statistics, optimization, probability theory, thermodynamics, economics, machine learning, to name a few.
[…]
blackOut of a rapidly growing literature we select two references that highlight the relevance of potential forces, i.e., generated as gradients of a potential, and hence to symmetric $A_{t}$’s. In the first one, Movilla Miangolarra etal. [33] study thermodynamic ensembles driven by potential forces to traverse a closed thermodynamic cycle, to obtain fundamental bounds on the maximal power that can be drawn from thermal anisotropy. In the next, Lavenant etal. [25] develop a framework of trajectory inference for cell differentiation and development, where the underlying potential forcing is due to the so-called Waddington’s landscape. In both situations, the constraint on vector fields governing the evolution of thermodynamic and cell states being potential arises from physical considerations.
[…]
Finally, a similar framework to the one developed herein should prove useful in ensemble control [28, 12, 7], where trajectories of multi-species populations indexed by $s$, for instance, are encoded in collections of transition matrices, $\Phi^{s}_{t}$’s, on a suitable bundle $\pi$ projecting onto the configuration in a base space, e.g., of covariance data of the form $\Sigma_{t}=\sum_{s}\Phi^{s}_{t}(\Phi^{s}_{t})^{\intercal}$. Other variants of practical interest include modeling the flow of particles based on aggregate information and the location of tracer particles used to estimate the holonomy of the collection, and, likewise, estimation of trajectories of dynamical systems driven by stochastic excitation based on sample trajectories. These are the subject of ongoing work.